The length of a rectangle is three times its width.
If the perimeter of the rectangle is 56m , find its area.

Answer by ReadingBoosters(3246)

p = 56 = 2l+2w
l = 3w
Substitute length, l, into the perimeter, p, equation
56 = 2(3w)+2w
Distribute the 2
56 = 6w + 2w
Combine like terms
56 = 8w
w = 7m is width
l = 3w = 3(7) = 21m is length
...
area = lxw = 7(21) = 147m^2
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.